Solve zx+zy = z2 with the initial conditions z(x, 0)=f(x)

1
Expert's answer
2021-01-18T06:17:09-0500

P=1

Q=1

R=z2

"\frac{dx}{P}=\frac{dy}{Q}=\frac{dz}{R}"

"\frac{dx}{1}=\frac{dy}{1}=\frac{dz}{z^2}"

Multiplyers: -z2,1,1

-z2dx+dy+dz=0

-z2x+y+z=C1

Multiplyers: 1,-z2,1

dx-z2dy+dz=0

x-z2y+z=C2

The general solution is,

"\phi(c_{1},c_{2})=0"

"\phi( -z^2x+y+z),(x-z^2y+z)=0"


-z2x+z+y-C1=0

D=1+4x(y-C1)

"z_1=\frac{-1+\sqrt{1+4x(y-C_1)}}{-2x}=\frac{-1+\sqrt{1-4xC_1}}{-2x}=f(x)"

"\sqrt{1-4xC_1}=-2xf(x)+1"

"4xC_1=-{(-2xf(x)+1)}^2+1"

"C_1(1)=\frac{-{(-2xf(x)+1)}^2+1}{4x}"

"z_2=\frac{-1-\sqrt{1+4x(y-C_1)}}{-2x}=\frac{-1-\sqrt{1-4xC_1}}{-2x}=f(x)"

"\sqrt{1-4xC_1}=2xf(x)-1"

"4xC_1=-{(2xf(x)-1)}^2+1"

"C_1(2)=\frac{-{(2xf(x)-1)}^2+1}{4x}"


-z2y+z+x-C2=0

As y=0

z=C2 - x= f(x)

C2=f(x) + x